OpenStudy (anonymous):

How do you solve for x if the equation is: x/3 = x+6/6

OpenStudy (rulnick):

if you mean x/3 = (x+6) / 6 then

OpenStudy (rulnick):

mult both sides by 6 to get 2x = x+6 subtract x from both sides to get x = 6

OpenStudy (anonymous):

thanks :)